OpenStudy (anonymous):

An object has a kinetic energy of 233 J and a momentum of magnitude 20.4 kg·m/s. Find the speed and the mass of the object.

OpenStudy (anonymous):

The equation for kinetic energy and the equation for momentum both have mass and velocity for variables. Setting each equal to the given amount, you have two equations and two unknowns. Solve for mass or velocity in one and plug it into the other to find one value. Use that value in the first equation to find the other. If you need more specifics, let me know.

OpenStudy (anonymous):

\[233=\frac{1}{2}mv^2=\frac{p^2}{2m}=\frac{20.4^2}{2m}\] Solve for m.

OpenStudy (anonymous):

Keaplurning can you work it out please.